# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ V.2 Exponential spatial complexity: Iterated Baumslag-Solitar
We now present an example of group dynamics that exhibits fragile fragmentation with a zero density jamming transition: the iterated Baumslag-Solitar group Clay and Margalit (2017). This group is (loosely speaking) constructed by embedding a ${\sf BS}$ group inside of itself. We refer to it as ${\sf BS}^{(2)}$, and define it via the presentation
@@
${\sf Dyn}_{{\sf BS}^{(2)}}$ dynamics are thus most naturally realized in spin-3 chains with 3-local dynamics. Note that ${\sf BS}^{(2)}$ also has a conserved $U(1)$ charge, this time given by the density of $\mathtt{c}$ generators, $n_{\mathtt{c}}\triangleq\sum_{i}n_{\mathtt{c},i}=\sum_{i}(|\mathtt{c}\rangle\langle\mathtt{c}|_{i}-|\mathtt{c}^{-1}\rangle\langle\mathtt{c}^{-1}|_{i})$. Several facts about ${\sf BS}^{(2)}$ (and related generalizations thereof) are proven in App. F. The most important result is that the Dehn function of ${\sf BS}^{(2)}$ is a super-exponential function of $L$:
@@
with the length of the RHS being doubly-exponentially larger than that of $|v(n)|$ (and where “$=$” in the above denotes equality as elements of ${\sf BS}^{(2)}$. By following the same strategy as in the construction of $w_{\rm large}(n)$, we can construct a word $w_{\rm huge}(n)\in K_{e,L}$ whose area grows doubly-exponentially with its length, viz.
@@
decreasing monotonically with time. The merit of taking this approach is that the irreversible setting allows us to extract lower bounds on the relaxation time of $n_{\mathtt{c}}$ for the reversible setting. Our simulations are run by initializing the system in $|w_{\rm huge}(n,L)\rangle$, a version of $w_{\rm huge}(n)$ padded with $L-|w_{\rm huge}(n)|$ $\mathtt{e}$ characters at random locations, so that $|w_{\rm huge}(n,L)|=L$, and then tracking the time evolution of $n_{|\mathtt{c}|}$. The results are shown in Fig. 10 for different values of $n$, which are necessarily very small on account of the doubly-exponential growth of the Dehn function. In the main panel, we show the relaxation time of $|w_{\rm huge}(3,L)\rangle$ as a function of $L$, defined by the time at which no $\mathtt{c},\mathtt{c}^{-1}$ characters remain in the evolved word. The inset shows ${\sf EL}(n)$, defined as the minimal value of $L$ for which relaxation (viz. the reaching of a state containing no $\mathtt{c},\mathtt{c}^{-1}$ characters) was observed to occur over 1000 runs of the dynamics. The extracted ${\sf EL}(n)$ roughly conforms to our expectation of ${\sf EL}(n)\sim 2^{\alpha n}$ for an $O(1)$ constant $\alpha$, although the long timescales required to observe thermalization mean that statistical errors are rather large, and with our current data we should not expect to obtain a perfectly exponential scaling.
@@ V.3.1 Defining fragile fragmentation
In other words, $\tau_{\mathcal{A}}$ is the ensemble of pure states one gets by evolving the initial state |$\ket{\psi}\otimes\ket{0}_{\mathcal{A}}$ for an arbitrary time, and post-selecting on the ancillas being in the final state |$\ket{0}_{\mathcal{A}}$. Note that in principle we could make $\tau_{\mathcal{A}}$ depend explicitly on the state of the ancillas. However, for group dynamics the most natural choice for this state is $|0\rangle_{\mathcal{A}}=|\mathtt{e}\rangle_{\mathcal{A}}$, and for semigroups an analogous state can be defined by augmenting the local Hilbert space with a character |$\ket{\mathtt{e}}$ which commutes with all characters. We therefore content ourselves with studying fragmentation for this particular choice of ancilla state. We define the Krylov sector of |$\ket{\psi}$ extended to $\mathcal{A}$ as
@@ V.3.2 Probing fragile fragmentation
Our definition of fragile fragmentation above makes reference to post-selection on the final state of the ancillas being |$\ket{0}_{\mathcal{A}}$. The probability of postselection succeeding is clearly exponentially small in $|\mathcal{A}|$. We now present a more efficient algorithm for (i) identifying whether a given system exhibits fragile fragmentation, and (ii) constructing the subspace $K_{\psi}(\mathcal{A})$ associated with an initial state |$\ket{\psi}$ given a maximum expansion length $L+|\mathcal{A}|$. This procedure is more efficient that naively postselecting on the state of the ancillas in various regimes which we discuss below. The general algorithm proceeds as follows. We start with the state |$\ket{\psi}\otimes\ket{0}_{\mathcal{A}}$, and evolve it under ${\sf Dyn}$ acting on the system plus ancillas for some time $t_{\mathrm{th}}$. After time $t_{\mathrm{th}}$, we repeat the following steps many times:
@@
The iteration stops when all ancilla sites have been decoupled: we know this is always possible since initial state was originally decoupled from the ancillas. On physical grounds we expect that the probability of getting outcome 00 in step 1 is $O(1)$ at all times, but for our purposes it suffices for it to scale as $1/{\rm poly}(L+|\mathcal{A}|)$. We first discuss how this algorithm can be used to construct the subspace $K_{\psi}(\mathcal{A})$. To accomplish this, one sets $t_{\mathrm{th}}$ to be the maximum possible thermalization time for the system plus ancillas, i.e., $t_{\mathrm{th}}\sim\exp(L+|\mathcal{A}|)$. Any fragmentation that persists after $t_{\mathrm{th}}$ will persist to infinite time for the given spatial resources. One can take the re-thermalization time after a measurement, $t_{\rm retherm}$, to be much shorter (i.e., as a low-order polynomial in $L+|\mathcal{A}|$), as the measurement is a single-site perturbation to the equilibrated state, and it is not expected to take more than polynomial time to have a significant amplitude to be in $|\mathtt{e}\rangle$. When the procedure terminates it yields a state |$\ket{\psi^{\prime}}$ that (by hypothesis) is in $K_{\psi}(\mathcal{A})$. After many runs, the ensemble of generated states spans $K_{\psi}(\mathcal{A})$. The procedure we described avoids the exponential overhead of post-selection, but still incurs the exponential overhead of mixing. If we want to reconstruct a state with overlap on all states in $K_{\text{in},\psi}$, this overhead cannot be avoided. Suppose, however, that we are not interested in full reconstruction of $K_{\text{in},\psi}$, but just in the simpler task of showing that adding ancillas and removing them (as above) partially lifts the fragmentation of the original system. More generally, suppose that we have constructed $K_{\psi}(\mathcal{A}_{0})$, and want to know if enlarging $\mathcal{A}_{0}$ to $\mathcal{A}_{1}$ enlarges the sector. We can run the initial equilibration step to a much shorter time than the full equilibration time. We stop when we have compressed back down to $\mathcal{A}_{0}$, and check if the resulting state is in $K_{\psi}(\mathcal{A}_{0})$  [^20] For word problems this can be done via a computational basis measurement; for systems with fragmentation in a non-product state basis, this measurement requires more care.. Finding a single state that lies outside $K_{\psi}(\mathcal{A}_{0})$ suffices to establish fragile fragmentation. Thus, detecting fragile fragmentation can be accomplished without requiring the system to fully thermalize or requiring a large bath. Can this procedure be used to study other properties of the fragmentation? One additional quantity that can be computed using this method are the geodesic lengths of group elements. Recall that the geodesic length $|g|$ of an element $g\in G$ is the length of the shortest word which represents $g$, viz. $|g|=\min\{|w|\,:\,\varphi(w)=g\}$. To compute this, we repeat this sequential length reducing procedure until the system freezes. Formally speaking, the system will freeze when $\matrixelement{\texttt{e}}{\rho_{\text{end}}}{\texttt{e}}=0$ for quantum dynamics (where $\rho_{\rm end}$ is the reduced density matrix of the spin at the end of the system), or $p(\texttt{e}_{\text{end}})=0$ for classical dynamics, where $p(\cdot)$ denotes a marginal distribution for the last site of the chain. When this occurs, the system size has reached the minimum length needed to support a word in the Krylov sector, therefore yielding the geodesic length of the word.
@@ V.3.3 Fragile fragmentation and thermalization
VI Beyond group dynamics
While we have focused on examples where the dynamics is constrained by a group multiplication law, the phenomena discussed so far are not limited to these models: as emphasized in Sec. II, the most general setting for word problem is that of semigroups. For semigroups, however, the geometric perspective adopted above is less useful [^22] To understand why, one may envision some notion of a Cayley 2-complex which labels elements of the semigroup. However, since there is no concept of an identity element, loops on the graph do not carry the same meaning that they did for groups. Furthermore, paths or loops on the Cayley 2-complex are no longer freely deformable because the semigroup relations are not invariant under cyclic conjugation, due to the lack of inverses. This holds true even for structures which are only mildly perturbed from groups, such as monoids.. In this section, we introduce two new semigroup models which do not admit a group structure but which nevertheless have word problems exhibiting large time and space complexity, which we establish using combinatorial rather than geometric arguments. The first example has large time and small space complexity, and shares similarity with the ${\sf BS}$ model. The second example has both large time and large space complexity, is qualitatively unique to non-group based dynamics, and leads to a more direct characterization of fragile fragmentation than provided by the general criterion of Sec. V.3.3. These models are inspired by the Motzkin spin chain and, more broadly, Motzkin dynamics (see Refs. Bravyi et al. (2012); Movassagh and Shor (2016)). For the readers’ convenience, we briefly summarize Motzkin dynamics. The local state space includes an identity character $|0\rangle$ (which plays the role of $|\mathtt{e}\rangle$ in our group-based models) as well as left and right parentheses |(,|)|($\rangle$,|$\rangle$| ( ⟩ , | ) ⟩. The dynamics is engineered so that the nestedness of the parenthesis remains preserved. For example, under the dynamics ‘()()( )’ may evolve to ‘()()()()( ) ( )’, but not ‘)()() (’.
@@ VI.1 Star-Motzkin model: slow thermalization
The allowed moves in the dynamics either create a pair of open/closed parentheses, or freely move a parenthesis past a ‘0’ character. To slow down the hydrodynamics of this conserved charge, we include an extra character $\ast$ in our Hilbert space, and modify the dynamics by adding the additional rules
[…]
i.e. when a parenthesis moves past a $\ast$ character (in a certain direction), the $\ast$ character is duplicated. The combination of these set of rules results in the dynamics which we will refer to as ${\sf Dyn}_{\ast M}$. One can readily see that ${\sf Dyn}_{\ast M}$ exhibits Hilbert space fragmentation. Indeed, if we ignore the $\ast$ character that was added, the Hamiltonian describes Motzkin dynamics, and thus already possesses fragmentation which cannot be described solely by a conserved parenthesis density. The sectors of these dynamics are labelled by a sequence of closed parentheses followed by a sequence of open parentheses taking the form $)^{m}(^{n}$: this corresponds to the total parenthesis imbalance in the configuration. When the $\ast$ character is added, a label for the Krylov sectors becomes
@@
Since $q$ can grow to be exponentially large in $L$, this last fact implies slow dynamics. In fact, one way to study this slow dynamics is to observe that atypical configurations of the $U(1)$ charge corresponding to the parentheses takes a very long time to thermalize. Regarding the second result, we also provide a crisper characterization for the circumstances under which it takes a long time to transition between two words (see G).
@@ VI.2 Chiral star-Motzkin model: fragile fragmentation
We now present an example where the expansion length is exponentially large, implying fragile fragmentation. In this example, we simply replace the $\ast$ character with a ‘chiral’ version of the character, which we denote as $\triangleright$. The rules for these new characters are similar to that of $\ast$, except that the rules are only activated when a $\triangleright$ character is adjacent to )))). More specifically, we replace the rules in (103) to
@@
VII Generalization to two dimensions: group loop models
The discussion thus far has been restricted to 1D models of group dynamics. It is natural to wonder whether or not higher-dimensional models with similar behavior can be constructed, especially since the aforementioned phenomena are more prevalent in higher dimensions. In this section, we discuss one 2D generalization of our group-based models that in some sense is the most faithful way of embedding the 1D group constraint in a two dimensional system, and which leads to a qualitatively new way of producing glassy dynamics and jamming in 2D. This proceeds by fixing a group $G$ (rather than a semigroup, but only for notational simplicity) and considering loop models that possess one flavor of loop for each generator of $G$. Along any one-dimensional reference loop, one can associate a group element corresponding to the product of all of the generators corresponding to loops that the reference loop intersects. The dynamics is engineered so that this group element remains invariant under the dynamics. This class of models can be viewed as a broad generalization of the construction in Refs. Stephen et al. (2022); Stahl et al. (2023) (which studied dynamics) and Ref. Balasubramanian et al. (2023) (which studied ground state properties), and we expect similar robustness of the Hilbert space fragmentation in these models. For brevity’s sake we will discuss this construction at a high level—e.g. we will largely use continuum language in order to avoid the notational burden incurred by an explicit lattice description—and will defer a more comprehensive analysis to future work. Given a discrete group $G$ and a presentation thereof, the degrees of freedom in our 2D model are associated with directed loops labeled by generators of $G$. In the following, we describe how to place constraints on the dynamics of these loops to produce phenomenology similar to that present in our 1D models. We start by observing that each directed reference path $\gamma$ through space (microscopically, along the lattice) can be associated with a group word $w(\gamma)$. This word is determined by the ordered labels of the loops that $\gamma$ intersects. Specifically, when proceeding along $\gamma$, each time $\gamma$ intersects a loop labeled by the generator $\mathtt{g}_{i}$, $w(\gamma)$ is i) multiplied by $\mathtt{g}_{i}$ if the local tangent vector $\vec{r}$ of $\gamma$ and the local tangent vector $\vec{\ell}$ of the loop can have a cross product $\vec{r}\times\vec{\ell}$ is parallel to ${\mathbf{\hat{z}}}$; and ii) is multiplied by $\mathtt{g}_{i}^{-1}$ if $\vec{r}\times\vec{\ell}$ is anti-parallel to ${\mathbf{\hat{z}}}$. This observation means that isolated closed loops can be regarded as implementing trivial relations in the words associated to the paths which pass through them, a fact we illustrate pictorially as

### 11 Apr 2024 v2

@@ V.2 Exponential spatial complexity: Iterated Baumslag-Solitar
We now present an example of group dynamics that exhibits fragile fragmentation with a zero density jamming transition: the iterated Baumslag-Solitar group [41]. This group is (loosely speaking) constructed by embedding a ${\sf BS}$ group inside of itself. We refer to it as ${\sf BS}^{(2)}$, and define it via the presentation
@@
${\sf Dyn}_{{\sf BS}^{(2)}}$ dynamics are thus most naturally realized in spin-3 chains with 3-local dynamics, whose local Hilbert space is obtained from that of the ${\sf BS}$ model by appending the two states $\{|\mathtt{c}\rangle,\,|\mathtt{c}^{-1}\rangle\}$. Note that like ${\sf BS}$, ${\sf BS}^{(2)}$ has a single conserved $U(1)$ charge given by the density of $\mathtt{c}$ generators, $n_{\mathtt{c}}\triangleq\sum_{i}n_{\mathtt{c},i}=\sum_{i}(|\mathtt{c}\rangle\langle\mathtt{c}|_{i}-|\mathtt{c}^{-1}\rangle\langle\mathtt{c}^{-1}|_{i})$. Several facts about ${\sf BS}^{(2)}$ (and related generalizations thereof) are proven in App. F. The most important result is that the Dehn function of ${\sf BS}^{(2)}$ is a super-exponential function of $L$:
@@
with the length of the RHS being doubly-exponentially larger than that of $|v(n)|$ (and where “$=$” in the above denotes equality as elements of ${\sf BS}^{(2)}$. By following the same strategy as in the construction of $w_{\rm large}(n)$, we can construct a word $w_{\rm huge}(n)\in K_{e,L}$ whose area grows doubly-exponentially with its length, namely
@@
decreasing monotonically with time. The merit of taking this approach is that the irreversible setting allows us to extract lower bounds on the relaxation time of $n_{\mathtt{c}}$ for the reversible setting. Our simulations are run by initializing the system in $|w_{\rm huge}(n,L)\rangle$, a version of $w_{\rm huge}(n)$ padded with $L-|w_{\rm huge}(n)|$ $\mathtt{e}$ characters at random locations, so that $|w_{\rm huge}(n,L)|=L$, and then tracking the time evolution of $n_{|\mathtt{c}|}$. The results are shown in Fig. 11 for different values of $n$, which are necessarily very small on account of the doubly-exponential growth of the Dehn function. In the main panel, we show the relaxation time of $|w_{\rm huge}(3,L)\rangle$ as a function of $L$, defined by the time at which no $\mathtt{c},\mathtt{c}^{-1}$ characters remain in the evolved word. The inset shows ${\sf EL}(n)$, defined as the minimal value of $L$ for which relaxation (namely the reaching of a state containing no $\mathtt{c},\mathtt{c}^{-1}$ characters) was observed to occur over 1000 runs of the dynamics. The extracted ${\sf EL}(n)$ roughly conforms to our expectation of ${\sf EL}(n)\sim 2^{\alpha n}$ for an $O(1)$ constant $\alpha$, although the long timescales required to observe thermalization mean that statistical errors are rather large, and with our current data we should not expect to obtain a perfectly exponential scaling.
@@ V.3.1 Defining fragile fragmentation
In other words, $\tau_{\mathcal{A}}$ is the ensemble of pure states one gets by evolving the initial state |$\ket{\psi}\otimes\ket{0}_{\mathcal{A}}$ for an arbitrary time, and postselecting on the ancillas being in the final state |$\ket{0}_{\mathcal{A}}$. Note that in principle we could make $\tau_{\mathcal{A}}$ depend explicitly on the state of the ancillas. However, for group dynamics the most natural choice for this state is $|0\rangle_{\mathcal{A}}=|\mathtt{e}\rangle_{\mathcal{A}}$, and for semigroups an analogous state can be defined by augmenting the local Hilbert space with a character |$\ket{\mathtt{e}}$ which commutes with all characters. We therefore content ourselves with studying fragmentation for this particular choice of ancilla state. We define the Krylov sector of |$\ket{\psi}$ extended to $\mathcal{A}$ as
@@ V.3.2 Probing fragile fragmentation
Our definition of fragile fragmentation above makes reference to postselection on the final state of the ancillas being |$\ket{0}_{\mathcal{A}}$. The probability of postselection succeeding is clearly exponentially small in $|\mathcal{A}|$. We now present a more efficient algorithm for (i) identifying whether a given system exhibits fragile fragmentation, and (ii) constructing the subspace $K_{\psi}(\mathcal{A})$ associated with an initial state |$\ket{\psi}$ given a maximum expansion length $L+|\mathcal{A}|$. This procedure is more efficient than naively postselecting on the state of the ancillas in various regimes which we discuss below. The general algorithm proceeds as follows. We start with the state |$\ket{\psi}\otimes\ket{0}_{\mathcal{A}}$, and evolve it under ${\sf Dyn}$ acting on the system plus ancillas for some time $t_{\mathrm{th}}$. After time $t_{\mathrm{th}}$, we repeat the following steps many times:
@@
The iteration stops when all ancilla sites have been decoupled: we know this is always possible since the initial state was originally decoupled from the ancillas. On physical grounds we expect that the probability of getting outcome 00 in step 1 is $O(1)$ at all times, but for our purposes it suffices for it to scale as $1/{\rm poly}(L+|\mathcal{A}|)$. We first discuss how this algorithm can be used to construct the subspace $K_{\psi}(\mathcal{A})$. To accomplish this, one sets $t_{\mathrm{th}}$ to be the maximum possible thermalization time for the system plus ancillas, i.e., $t_{\mathrm{th}}\sim\exp(L+|\mathcal{A}|)$. Any fragmentation that persists after $t_{\mathrm{th}}$ will persist to infinite time for the given spatial resources. One can take the re-thermalization time after a measurement, $t_{\rm retherm}$, to be much shorter (i.e., as a low-order polynomial in $L+|\mathcal{A}|$), as the measurement is a single-site perturbation to the equilibrated state, and it is not expected to take more than polynomial time to have a significant amplitude to be in $|\mathtt{e}\rangle$. When the procedure terminates it yields a state |$\ket{\psi^{\prime}}$ that (by hypothesis) is in $K_{\psi}(\mathcal{A})$. After many runs, the ensemble of generated states spans $K_{\psi}(\mathcal{A})$. The procedure we described avoids the exponential overhead of postselection, but still incurs the exponential overhead of mixing. If we want to reconstruct a state with overlap on all states in $K_{\text{in},\psi}$, this overhead cannot be avoided. Suppose, however, that we are not interested in full reconstruction of $K_{\text{in},\psi}$, but just in the simpler task of showing that adding ancillas and removing them (as above) partially lifts the fragmentation of the original system. More generally, suppose that we have constructed $K_{\psi}(\mathcal{A}_{0})$, and want to know if enlarging $\mathcal{A}_{0}$ to $\mathcal{A}_{1}$ enlarges the sector. We can run the initial equilibration step to a much shorter time than the full equilibration time. We stop when we have compressed back down to $\mathcal{A}_{0}$, and check if the resulting state is in $K_{\psi}(\mathcal{A}_{0})$  [^21] For word problems this can be done via a computational basis measurement; for systems with fragmentation in a non-product state basis, this measurement requires more care.. Finding a single state that lies outside $K_{\psi}(\mathcal{A}_{0})$ suffices to establish fragile fragmentation. Thus, detecting fragile fragmentation can be accomplished without requiring the system to fully thermalize or requiring a large bath. Can this procedure be used to study other properties of the fragmentation? One additional quantity that can be computed using this method are the geodesic lengths of group elements. Recall that the geodesic length $|g|$ of an element $g\in G$ is the length of the shortest word which represents $g$, namely $|g|=\min\{|w|\,:\,\varphi(|w\rangle)=g\}$. To compute this, we repeat this sequential length reducing procedure until the system freezes. Formally speaking, the system will freeze when $\matrixelement{\texttt{e}}{\rho_{\text{end}}}{\texttt{e}}=0$ for quantum dynamics (where $\rho_{\rm end}$ is the reduced density matrix of the spin at the end of the system), or $p(\texttt{e}_{\text{end}})=0$ for classical dynamics, where $p(\cdot)$ denotes a marginal distribution for the last site of the chain. When this occurs, the system size has reached the minimum length needed to support a word in the Krylov sector, therefore yielding the geodesic length of the word.
@@ V.3.3 Fragile fragmentation and thermalization
VI Semigroup examples
In the explicit examples of ${\sf Dyn}_{G}$ dynamics studied above, $G$ has been taken to have the structure of a group. The phenomena discussed so far are however not limited to models with a group structure; indeed all of the general results obtained in Secs. II and III were are valid for any finitely presented semigroup. For semigroups, however, the geometric perspective adopted above in the discussion of ${\sf Dyn}_{\sf BS}$ is less useful [^23] To understand why, one may envision some notion of a Cayley 2-complex which labels elements of the semigroup. However, since there is no concept of an identity element, loops on the graph do not carry the same meaning that they do for groups. Furthermore, paths or loops on the Cayley 2-complex are no longer freely deformable because the semigroup relations are not invariant under cyclic conjugation, due to the lack of inverses.. In this section, we introduce two new semigroup models which do not admit a group structure but which nevertheless have word problems exhibiting large time and space complexity, which we establish using combinatorial rather than geometric arguments. The first example has large time and small space complexity, and shares similarity with the ${\sf BS}$ model. The second example has both large time and large space complexity, is qualitatively unique to non-group based dynamics, and leads to a more direct characterization of fragile fragmentation than that provided by the general criterion of Sec. V.3.3. These models are inspired by the Motzkin spin chain and, more broadly, Motzkin dynamics (see Refs. [51, 58]). For the readers’ convenience, we briefly summarize Motzkin dynamics. The local state space includes an identity character $|0\rangle$ (which plays the role of $|\mathtt{e}\rangle$ in our group-based models) as well as left and right parentheses |(,|)|($\rangle$,|$\rangle$| ( ⟩ , | ) ⟩. The dynamics is engineered so that the “nestedness” of the parenthesis remains preserved, where nestedness is defined by the number of left parenthesis located to the left of matching right parenthesis; for example, under the dynamics the word ‘()()( )’ may evolve to ‘()()()()( ) ( )’, ‘(())(())( ( ) )’ or ‘00’, but not to ‘)()() (’, ‘((((( (’, or ‘))))) )’. These rules can be summarized formally by defining the Motzkin semigroup as
${\sf Motz}={\sf semi}\langle 0,(,)\,|\,(0=0(,\,)0=0),\,()=00\rangle.$ (106)
Alternatively, we may define a height field $h_{i}$ which keeps track of the level of nestedness at site $i$, via
$h_{i}=\sum_{j<i}(|(\rangle\langle(|_{j}-|)\rangle\langle)|_{j}).$ (107)
The dynamics then preserves both $h_{L}$ (the net difference between the number of (((( and )))) parenthesis) and $\min_{i}h_{i}$ (which measures the extent of the nestedness).
@@ VI.1 Star-Motzkin model: slow thermalization
Thus when a parenthesis moves past a $\ast$ character (in a certain direction), the $\ast$ character is duplicated. The combination of these set of rules results in the dynamics which we will refer to as ${\sf Dyn}_{\ast M}$. One can readily see that ${\sf Dyn}_{\ast M}$ exhibits Hilbert space fragmentation. Indeed, if we ignore the $\ast$ character that was added, the Hamiltonian describes Motzkin dynamics, and thus already possesses fragmentation which cannot be described solely by a conserved parenthesis density. The sectors of these dynamics are labelled by a sequence of closed parentheses followed by a sequence of open parentheses taking the form $)^{m}(^{n}$: this corresponds to the total parenthesis imbalance in the configuration. When the $\ast$ character is added, a label for the Krylov sectors becomes
@@
Since $q$ can grow to be exponentially large in $L$, this last fact implies slow dynamics. In fact, one way to study this slow dynamics is to observe that atypical configurations of the $U(1)$ charge corresponding to the parentheses takes a very long time to thermalize. Regarding the second result, we also provide a crisper characterization for the circumstances under which it takes a long time to transition between two words (see App. G).
@@ VI.2 Chiral star-Motzkin model: fragile fragmentation
We now present an example where the expansion length is exponentially large, implying fragile fragmentation. In this example, we simply replace the $\ast$ character with a ‘chiral’ version of the character, which we denote as $\triangleright$. The rules for these new characters are similar to that of $\ast$, except that the rules are only activated when a $\triangleright$ character is adjacent to )))). More specifically, we replace the rules in (108) with
@@
VII Generalizations to two dimensions: group loop models
The discussion thus far has been restricted to 1D models of group dynamics. It is natural to wonder whether or not higher-dimensional models with similar behavior can be constructed, especially since the aforementioned phenomena are more prevalent in higher dimensions. In this section, we discuss one 2D generalization of our group-based models that in some sense is the most faithful way of embedding the 1D group constraint in a two dimensional system, and which leads to a qualitatively new way of producing glassy dynamics and jamming in 2D. This proceeds by fixing a group $G$ [^24] As we will see shortly, our 2D models are simplest to write down when $G$ is a group rather than a semigroup, although this is not a fundamental restriction. and considering loop models that possess one flavor of loop for each generator of $G$. Along any one-dimensional reference loop, one can associate a group element corresponding to the product of all of the generators corresponding to loops that the reference loop intersects. The dynamics is engineered so that this group element remains invariant under the dynamics. This class of models can be viewed as a broad generalization of the construction in Refs. [59, 60] (which studied dynamics) and Ref. [61] (which studied ground state properties), and we expect similar robustness of the Hilbert space fragmentation in these models. For brevity’s sake we will discuss this construction at a high level—e.g. we will largely use continuum language in order to avoid the notational burden incurred by an explicit lattice description—and will defer a more comprehensive analysis to future work. Given a discrete group $G$ and a presentation thereof, the degrees of freedom in our 2D model are associated with directed loops labeled by generators of $G$. In the following, we describe how to place constraints on the dynamics of these loops to produce phenomenology similar to that present in our 1D models. We start by observing that each directed reference path $\gamma$ through space (microscopically, along the lattice) can be associated with a group word $w(\gamma)$. This word is determined by the ordered labels of the loops that $\gamma$ intersects. Specifically, when proceeding along $\gamma$, each time $\gamma$ intersects a loop labeled by the generator $\mathtt{g}_{i}$, $w(\gamma)$ is i) multiplied by $\mathtt{g}_{i}$ if the local tangent vector $\vec{r}$ of $\gamma$ and the local tangent vector $\vec{\ell}$ of the loop can have a cross product $\vec{r}\times\vec{\ell}$ is parallel to ${\mathbf{\hat{z}}}$; and ii) is multiplied by $\mathtt{g}_{i}^{-1}$ if $\vec{r}\times\vec{\ell}$ is anti-parallel to ${\mathbf{\hat{z}}}$. This observation means that isolated closed loops can be regarded as implementing trivial relations in the words associated to the paths which pass through them, a fact we illustrate pictorially as